# Emergence of two inertial sub-ranges in solar wind turbulence: dependence on heliospheric distance and solar activity

### v1

@@ Abstract
The solar wind is highly turbulent, and intermittency effects are observed for fluctuations within the inertial range. By analyzing magnetic field spectra and fourth-order moments, we perform a comparative study of intermittency in different types of solar wind measured during periods of solar minima and a maximum. Using eight fast solar wind intervals measured during solar minima between 0.3 au and 3.16 au, we found a clear signature of two inertial sub-ranges with $f^{-3/2}$ and $f^{-5/3}$ power laws in the magnetic power spectra. The intermittency, measured through the scaling law of the kurtosis of magnetic field fluctuations, further confirms the existence of two different power laws separated by a clear break. A systematic study on the evolution of the said sub-ranges as a function of heliospheric distance shows correlation of the break scale with both the turbulence outer scale and the typical ion scales. During solar maximum, we analyzed five intervals for each of Alfvénic fast, Alfvénic slow and non-Alfvénic slow solar wind. Unlike the case during the solar minima, the two sub-ranges are no longer prominent and the Alfvénic slow wind is found to be in an intermediate state of turbulence compared to that of the fast wind and the usual non-Alfvénic slow wind.
@@ 1 Introduction
The solar wind is the most accessible natural laboratory for studying turbulence in space plasmas (Bruno & Carbone, 2013). The dynamic solar activity and the diversity of the originating regions produce solar wind with a variety of characteristics, the most evident being the plasma speed. While the fast solar wind (FSW, $>550\ km\ s^{-1}$) mainly emanates from the polar coronal holes, the slow solar wind (SSW, $<400\ km\ s^{-1}$) is believed to be originated from equatorial streamers (Belcher & Davis Jr., 1971; Smith et al., 1978; Phillips et al., 1995). During high solar activity, however, both FSW and SSW are distributed at all latitudes instead of being confined exclusively to polar and equatorial regions, respectively. Another interesting feature of FSW is the high Alfvénicity $i.e.$ high correlation (or anti-correlation) between velocity fluctuations (v) and b ($=\mathbf{B}/\sqrt{\mu_{0}\rho}$, where B the magnetic field fluctuation and $\rho$ the mass density) in contrast with the SSW comprising of weak ${\bf v}$-${\bf b}$ correlations. This one-to-one correspondence, however, does not strictly hold during high solar activity, as a third type of wind is also observed. This wind, termed as Alfvénic slow solar wind (ASSW), has speed similar to that of the slow wind but is surprisingly permeated with high Alfvénicity (Marsch et al., 1981; D’Amicis et al., 2011; D’Amicis & Bruno, 2015). The degree of Alfvénicity influences the nature of turbulence in different types of solar winds. A high degree of Alfvénicity represents an imbalance between the Elsässer variables, $\mathbf{z^{\pm}}=\textbf{v}\pm\textbf{b}$, thus leading to less developed turbulence whereas low Alfvénicity corresponds to comparatively more developed turbulence owing to the balance between them. At scales greater than the ion-inertial length ($d_{i}$), a longer $k^{-5/3}$ energy power spectrum is therefore observed for the slow wind whereas a comparatively shorter $k^{-5/3}$ spectrum is observed in the Alfvénic fast wind (Bruno & Carbone, 2005, 2013; D’Amicis et al., 2018).
These observed spectra are universal in solar wind turbulence and are consistent with self-similar energy cascade (Kolmogorov phenomenology) within the inertial range. In physical space, universal energy cascade is obtained in terms of the linear scaling law for the third-order moments of velocity and magnetic field fluctuations (Sorriso-Valvo et al., 2007; Banerjee et al., 2016; Marino & Sorriso-Valvo, 2023). In order to assure a self-similar cascade, the kurtosis $K$ (the normalised fourth-order moment of the fluctuations) should be scale invariant. For a turbulent flow, one such possibility is the case of quasi-Gaussian PDFs where the third-order moment (skewness) is non-zero but the K is roughly equal to that of a Gaussian distribution.
However, careful studies in turbulent fluids and plasmas consistently show a departure from self-similarity as one moves towards the smaller length scales within the inertial range. This departure, known as inertial-range intermittency, is characterised by the large tails of the PDFs at those scales. In particular, intermittency effects are quantified by the deviation from the self-similar scaling laws of the higher-order moments (Frisch, 1995; Biskamp, 2003; Banerjee, 2014). Instead of using arbitrary higher-order moments, the kurtosis is often used as a practical measure of intermittency and a higher probability of extreme events leads to its increase with decreasing length scale $\ell$. From a physical point of view, this implies that the small-scale coherent structures, such as the vortices, current sheets, etc., generated due to nonlinear interactions, do not fill the available space in a self-similar way nor are randomly distributed, but rather tend to form inhomogeneously distributed clusters of bursts (Frisch, 1995; Sorriso-Valvo et al., 1999; Bruno et al., 2003).
@@
While the solar wind expands and accelerates through the heliosphere, the turbulence becomes more developed, with the fluctuations being majorly energized by the nonlinear interactions between the oppositely propagating Alfvén waves (Chandran, 2018), switchbacks (Bale et al., 2021; Sakshee et al., 2022), large-scale structures, and instabilities (Bavassano et al., 1982a; Roberts et al., 1992). Studies based on spacecraft observations have shown a variation in spectral indices of the magnetic and velocity power spectra (Bavassano et al., 1982b; Roberts, 2010; Chen et al., 2020), a decrease in v-b correlations and a broadening of the inertial range (Bavassano et al., 1998, 1982b; Davis et al., 2023). Recently, using high resolution in-situ data of the Parker Solar Probe it has been suggested that the magnetic spectral index evolves from $-3/2$ near the Sun (as close as 0.17 au) to a more developed $-5/3$ at 1 au (Alberti et al., 2020; Chen et al., 2020; Shi, C. et al., 2021; Sioulas et al., 2023a). These observations are consistent with the idea of radial evolution of solar wind turbulence into more developed states and the non-adiabatic heating of the medium with increasing heliospheric distance (Marsch et al., 1982; Cranmer et al., 2009; Hellinger et al., 2011). In addition, a power law behaviour for the kurtosis (Bruno et al., 2003; Di Mare et al., 2019; Carbone et al., 2021; Hernández et al., 2021) and an increase in intermittency in solar wind turbulence have been observed at greater heliospheric distances (Sioulas et al., 2022; Sorriso-Valvo et al., 2023). Using the magnetic data of Helios 2, Sorriso-Valvo et al. (2023) observed a break in the scaling of $K$ of the magnetic field fluctuations in FSW, during solar minimum. They provided a plausible explanation suggesting this observed break to be associated with the $f^{-1}$ break in the magnetic power spectrum of FSW. However, clear disparity is observed between the scales corresponding to the breaks in kurtosis scaling and the power spectra. A break in both the spectral density and the higher-order structure functions has also been observed (Wicks et al., 2011; Wu et al., 2022; Telloni, 2022; Sioulas et al., 2023b; Wu et al., 2023). However, the nature of such break and its implications on the dynamics of the solar wind turbulence have not been investigated in detail yet.
@@
In this paper, we revisit the aforementioned problem and carry out a systematic study to provide an explanation for the break observed in the kurtosis scaling. Using the in-situ data of Helios and Ulysses during solar minima, we show the kurtosis break is primarily associated with an observed break between two inertial sub-regimes of magnetic power spectra, having $-3/2$ and $-5/3$ spectral indices, respectively. In addition, we also study the radial evolution of the break scale to characterise the solar wind turbulence as a function of the heliospheric distance. During a solar maximum, however, breaks are not prominent in the scaling of K. Nevertheless, a comparative study of FSW, ASSW and SSW shows a clear distinction of these three types of wind according to the degree of turbulence and the degree of intermittency, along with some insights on the origin of ASSW. In Sections 2 and 3, we briefly describe the data and methodologies used for the analysis. Section 4 provides the results obtained in our study during solar minima (4.1) and maxima (4.2), respectively. Finally, in Section 5, we summarize our findings and conclude.
@@ 2 DATA SELECTION
Label Year Time Interval $v_{sw}$ R
(MM-DD-HH) (km/sec) (au)
@@
Figure 1: FSW intervals indicated by red boxes using Helios (left) and Ulysses (right) data, during solar minima. Top to bottom: proton number density, solar wind speed, interplanetary magnetic field, correlation co-efficient between the components of proton velocity and magnetic field computed over a 10 hr window. Time (x) axis is in the format: MM:DD:HH.
@@
Figure 2: Different Ulysses intervals of SW during a period of solar maximum. Top to bottom: proton number density, solar wind speed, interplanetary magnetic field, correlation co-efficient between the components of proton velocity and magnetic field computed over a 10 hr window. Time (x) axis is in the format: MM:DD:HH. Blue, green and red boxes represent SSW, ASSW, and FSW intervals, respectively.
@@
For our analysis, we have used in-situ data from the Helios and Ulysses spacecraft data repository publicly available at NASA CDAWeb (https://cdaweb.gsfc.nasa.gov) and AMDA science analysis system (https://amda.irap.omp.eu). The plasma data for Helios and Ulysses have been obtained from the E1 Plasma Experiment instrument and the Solar Wind Observations Over the Poles of the Sun (SWOOPS) instrument, respectively. For magnetic power spectrum and kurtosis scaling, we use 6 s resolution magnetic-field data from the E3 Flux-gate Magnetometer (FGM) onboard Helios and 1 s resolution magnetic field data from the Vector Helium Magnetometer (VHM) onboard Ulysses spacecraft. During a declining phase of solar activity near a solar minimum between 1975 and 1976, Helios 1 and 2 recorded several streams of FSW from a coronal hole (or the same source), which sustained through nearly two solar rotations (Bruno et al., 2003). Several intervals of the fast wind expelled from this coronal hole were also identified by Perrone et al. (2018). In particular, for our current analysis, we use the streams - A3, A6, A7, and A8, ranging from 0.3 au to 1 au, mentioned therein. Each chosen interval (i) contains negligibly small amount of data gaps, (ii) is free of any considerable mean trend, and (iii) turns up to be reasonably stationary. The stationarity is assured by the approximate constant average of sub-intervals of different lengths. Extending our analysis beyond 1 au, we use four intervals of FSW at varying heliospheric distances (F1 - F4 as listed in Table - 1), recorded by Ulysses during the years 1995-1996. Typical features of certain FSW intervals in the inner and outer heliosphere used in our analysis with high v-b correlations are shown in Fig. 1.
@@
In order to interpret our findings, we also need to compute the co-spectra of cross-helicty $\sigma_{c}$ (see Section 3), for which we have used the 40.5 s resolution magnetic field and proton velocity data from the E3 FGM and the E1 Plasma Experiment instrument onboard Helios. We use degraded resolution for the magnetic field data in order to keep coherence with the available plasma data from the data repository. A similar analysis cannot be done using the plasma data of Ulysses where the data resolution is 240 s, and hence cannot be used to capture the required length scales of our interest.
@@ 3 ANALYSIS METHOD
Our analysis is mainly based on the computation of (i) the kurtosis ($K$) or the normalized fourth-order moments of magnetic field fluctuations, (ii) the magnetic power spectral density ($PSD$), and (iii) the cross-helicity co-spectra ($\hat{\sigma_{c}}$). All the data sets were made evenly sampled by interpolating the data gaps before using for any of the computations.
Since all the intervals used in our study are permeated by superalfvénic solar wind, one can practically use Taylor’s hypothesis, which means if the turbulent fluctuations are much smaller than the bulk speed, they can be considered as frozen (or slowly evolving) as the flow sweeps the probe (Taylor, 1938). When using single-point measurements in the form of a time series, the only accessible direction for the increments is along the bulk flow. This provides an equivalence between the longitudinal (along the flow) length scale $\ell$ and the corresponding time scale $\tau$ as ${\ell}=V_{sw}\tau$, where $V_{sw}$ is the mean solar wind speed. Therefore, we define the increments of the $i^{th}$ component (with $i=r,t,n$) of the magnetic field as $\Delta b_{i}(t,\tau)=b_{i}(t+\tau)-b_{i}(t)$. In order to capture both magnitudinal and directional fluctuations of b , we define the $n^{th}$ order structure function as:
@@
Note that, when each $\Delta b_{i}$ follows a Gaussian distribution with zero mean, $K(\tau)$ is equal to $5/3$ (see appendix Section A). For a self-similar, non intermittent flow, in the inertial range of scales (namely much smaller than the energy-injection scales and larger than the dissipative scales) the $n^{th}$ order structure function is expected to scale as $S_{n}(\tau)\propto\tau^{np}$, where $p$ is a phenomenological constant (Frisch, 1995). It is therefore straightforward to see that $K$ becomes independent of $\tau$. However, in the presence of intermittency, this linear scaling does not hold any longer and the simplest intermittency model can be given as $S_{n}(\tau)\propto\tau^{np+q{(n)}}$, where $q{(n)}$ is a nonlinear correction accounting for the intermittent structures. For the kurtosis, this leads to a power-law scaling $K(\tau)\sim\tau^{-\kappa}$, with $\kappa=q{(4)}/2q{(2)}$. Such a scaling, universally observed in fluid turbulence, has recently been described in the case of solar wind turbulence as well (Di Mare et al., 2019; Hernández et al., 2021; Sorriso-Valvo et al., 2023). In this work, we study the scaling properties of $K$ of the magnetic field fluctuations at different heliospheric distances.
Finally, the magnetic energy spectra and cross-helicity co-spectra are defined by, $PSD=\hat{b_{i}}^{\dagger}\hat{b_{i}}$ and $\hat{\sigma_{c}}=(\hat{b_{i}}^{\dagger}\hat{v_{i}}+\hat{v_{i}}^{\dagger}\hat{b_{i}})/2$ respectively, where $\hat{b_{i}}$ and $\hat{v_{i}}$ are the Fast Fourier transforms (FFT) of the magnetic and velocity field components $b_{i}$ and $v_{i}$, respectively, with summation being intended over the repeated indices (where $i=r,t,n$).

### v2

@@ Abstract
The solar wind is highly turbulent, and intermittency effects are observed for fluctuations within the inertial range. By analyzing magnetic field spectra and fourth-order moments, we perform a comparative study of turbulence and intermittency in different types of solar wind measured during periods of solar minima and a maximum. Using eight fast solar wind intervals measured during solar minima between 0.3 au and 3.16 au, we found a clear signature of two inertial sub-ranges with $f^{-3/2}$ and $f^{-5/3}$ power laws in the magnetic power spectra. The intermittency, measured through the scaling law of the kurtosis of magnetic field fluctuations, further confirms the existence of two different power laws separated by a clear break. A systematic study on the evolution of the said sub-ranges as a function of heliospheric distance shows correlation of the break scale with both the turbulence outer scale and the typical ion scales. During solar maximum, on the contrary, the two sub-ranges are not omnipresent, thus showing more variability in the power spectra and intermittency scaling properties.
@@ 1 Introduction
The highly dynamic solar activity and the diversity of the originating regions produce solar wind with a variety of characteristics, the most evident being the plasma speed. While the fast solar wind (FSW, $>550\ km\ s^{-1}$) mainly emanates from the polar coronal holes, the slow solar wind (SSW, $<400\ km\ s^{-1}$) is believed to be originated from equatorial streamers (Belcher & Davis Jr., 1971; Smith et al., 1978; Phillips et al., 1995). During high solar activity, however, both FSW and SSW are distributed at all latitudes instead of being confined exclusively to polar and equatorial regions, respectively. An interesting feature of FSW is the high Alfvénicity, namely the high correlation (or anti-correlation) between fluctuations in velocity and magnetic field $\textbf{b}=\mathbf{B}/\sqrt{\mu_{0}\rho}$ (where the normalization using the mass density $\rho$ is used to convert magnetic field into velocity units). In contrast, the SSW predominantly comprises of weak ${\bf v}$-${\bf b}$ correlations. This one-to-one correspondence, however, does not strictly hold during high solar activity, as a third type of wind is also observed. This wind, termed as Alfvénic slow solar wind (ASSW), has low speed but is surprisingly permeated with high Alfvénicity (Marsch et al., 1981; D’Amicis et al., 2011; D’Amicis & Bruno, 2015). The degree of Alfvénicity correlates with the nature of turbulence in different types of solar winds (see Bruno & Carbone, 2013, and references therein). A high degree of Alfvénicity corresponds to an imbalance between the inward and outward Alfvénic fluctuations of solar origin that propagate along the heliospheric magnetic field, which, according to MHD models, results in reduced nonlinear interactions and thus less developed turbulence (Kraichnan, 1965; Bruno & Carbone, 2013). On the contrary, low Alfvénicity corresponds to comparatively more developed turbulence, owing to the stronger nonlinear interactions between balanced inward and outward perturbations. In solar wind streams, a broader $k^{-5/3}$ energy power spectrum is therefore observed for the slow wind, whereas a comparatively shorter spectrum is observed in the Alfvénic fast and slow winds (Bruno & Carbone, 2005, 2013; D’Amicis et al., 2018, 2022).
@@
While the solar wind expands and accelerates through the heliosphere, the turbulence becomes more developed, with the fluctuations being majorly energized by the nonlinear interactions between the oppositely propagating Alfvén fluctuations (Chandran, 2018), switchbacks (Bale et al., 2021; Sakshee et al., 2022), large-scale structures, and instabilities (Bavassano et al., 1982a; Roberts et al., 1992). Using in-situ spacecraft data, a steepening of the magnetic power spectra was found with increasing heliospheric distance, $R$, in the inner-heliosphere and beyond $1$ au (Bavassano et al., 1982b; Roberts, 2010). In addition, a decrease in v-b correlations and a broadening of the inertial range was also found as $R$ increases (Bavassano et al., 1998, 1982b; Davis et al., 2023). Recently, using high resolution data of the Parker Solar Probe it has been suggested that the magnetic spectral index evolves from a shallower $-3/2$ near the Sun (as close as 0.17 au), typical of strongly Alfvénic MHD turbulence to a more developed $-5/3$ at 1 au (Alberti et al., 2020; Chen et al., 2020; Shi, C. et al., 2021; Sioulas et al., 2023a). These observations are consistent with the idea of radial evolution of solar wind turbulence into more developed states and the corresponding non-adiabatic heating of the medium with increasing heliospheric distance (Marsch et al., 1982; Cranmer et al., 2009; Hellinger et al., 2011). In addition, a power law behaviour for the kurtosis (Bruno et al., 2003; Di Mare et al., 2019; Carbone et al., 2021; Hernández et al., 2021) and an increase in intermittency in solar wind turbulence have been observed at greater heliospheric distances (Sioulas et al., 2022; Sorriso-Valvo et al., 2023).
Recently, more careful analysis of scaling laws are emerging that challenge this well established framework. Preliminary studies suggest that a break in both the spectral density and the higher-order structure functions seems to characterize strongly Alfvénic solar wind intervals (Wicks et al., 2011; Wu et al., 2022; Telloni, 2022; Sioulas et al., 2023b; Wu et al., 2023; Sorriso-Valvo et al., 2023). This would introduce a substantial modification of the scaling invariance of the solar wind MHD fluctuations. However, the nature of such break and its implications on the dynamics of the solar wind turbulence have not been investigated in detail yet.
@@
In this paper, we revisit the aforementioned problem and carry out a systematic study to provide a detailed description of the break observed in the solar wind scaling laws of turbulent fluctuations. Using in-situ data of Helios and Ulysses during solar minima, we show that the break observed in spectra and kurtosis consistently separate two inertial sub-regimes, having $-3/2$ and $-5/3$ spectral indices, and weaker and stronger intermittency, respectively. In addition, we also study the radial evolution of the break scale to characterise the solar wind turbulence as a function of the heliospheric distance. During a solar maximum, a comparative study of FSW, ASSW and SSW measured by Ulysses shows a less clear classification of the inertial range, with breaks still being mostly present in Alfvénic streams. In Sections 2 and 3, we briefly describe the data and methodologies used for the analysis. Section 4 provides the results obtained in our study during solar minima (4.1) and maxima (4.2), respectively. Finally, in Section 5, we summarize our findings and conclude.
@@ 2 DATA SELECTION
Label Year Time Interval $V_{sw}$ R lat
(MM-DD-HH) (km/sec) (au) (°$\degree$)
@@
Figure 1: FSW intervals indicated by red boxes using Helios (left) and Ulysses (right) data, during solar minima. Top to bottom: proton number density, solar wind speed, interplanetary magnetic field, correlation coefficient between the components of proton velocity and magnetic field computed over a 2 hr window.
@@
Figure 2: Different Ulysses intervals of SW during a period of solar maximum. Top to bottom: proton number density, solar wind speed, interplanetary magnetic field, correlation co-efficient between the components of proton velocity and magnetic field computed over a 2 hr window. Blue, green and red boxes represent SSW, ASSW, and FSW intervals, respectively.
@@
For our analysis, we have used in-situ data from the Helios and Ulysses spacecraft data repository publicly available at NASA CDAWeb (https://cdaweb.gsfc.nasa.gov) and AMDA science analysis system (https://amda.irap.omp.eu). The plasma data for Helios and Ulysses have been obtained from the E1 Plasma Experiment instrument and the Solar Wind Observations Over the Poles of the Sun (SWOOPS) instrument, respectively. For magnetic power spectrum and kurtosis scaling, we use 6 s resolution magnetic-field data from the E3 Flux-gate Magnetometer (FGM) onboard Helios and 1 s resolution magnetic field data from the Vector Helium Magnetometer (VHM) onboard Ulysses spacecraft. During a declining phase of solar activity near a solar minimum between 1975 and 1976, Helios 1 and 2 recorded several streams of FSW from a coronal hole (or the same source), which sustained through nearly two solar rotations (Bruno et al., 2003). Several intervals of the fast wind expelled from this coronal hole were also identified by Perrone et al. (2018). In particular, for our current analysis, we use the streams A3, A6, A7, and A8, ranging from 0.3 au to 1 au, mentioned therein. Each chosen interval contains negligibly small amount of data gaps, is free of any considerable mean trend, and turns up to be reasonably stationary. The stationarity is assured by the approximate constant average of sub-intervals of different lengths. Extending our analysis beyond 1 au, we use four intervals of FSW at varying heliospheric distances (F1—F4 as listed in Table 1), recorded by Ulysses during the years 1995-1996. Typical features of certain FSW intervals in the inner and outer heliosphere used in our analysis with high ${\bf v}$-${\bf b}$ correlations are shown in Fig. 1.
@@
In order to interpret our findings, we also need to compute the co-spectra of cross-helicity $\sigma_{c}$ (see Section 3), for which we have used the 40.5 s resolution magnetic field and proton velocity data from the E3 FGM and the E1 Plasma Experiment instrument onboard Helios. We use degraded resolution for the magnetic field data in order to keep coherence with the available plasma data from the data repository. A similar analysis cannot be done using the plasma data of Ulysses where the data resolution is 240 s, and hence cannot be used to capture the required length scales of our interest.
@@ 3 ANALYSIS METHOD
Our analysis is mainly based on the computation of (i) the magnetic power spectral density ($PSD$), (ii) the kurtosis ($K$) or the normalized fourth-order moments of magnetic field fluctuations, and (iii) the cross-helicity co-spectra ($\hat{\sigma_{c}}$). All of the data sets were made evenly sampled by interpolating the data gaps before any of the computations.
Since all the intervals used in our study are permeated by super-Alfvénic solar wind, one can practically use Taylor’s hypothesis, which means if the phase speed of the fluctuations is much smaller than the flow speed of the solar wind, the fluctuations can be considered as frozen (or slowly evolving) as the flow sweeps the probe (Taylor, 1938). When using single-point measurements in the form of a time series, the only accessible direction for the increments is along the bulk flow. This provides an equivalence between the longitudinal (along the flow) length scale $\ell$ and the corresponding time scale $\tau$ as ${\ell}=V_{sw}\tau$, where $V_{sw}$ is the mean solar wind speed. Therefore, we define the increments of the $i^{th}$ component (with $i=r,t,n$ indicating the vector components in the standard RTN coordinate system) of the magnetic field as $\Delta B_{i}(t,\tau)=B_{i}(t+\tau)-B_{i}(t)$. In order to capture both magnitudinal and directional fluctuations of $\mathbf{B}$, we define the $n^{th}$ order structure function as:
@@
Note that, when each $\Delta B_{i}$ follows a Gaussian distribution with zero mean, $K(\tau)$ is equal to $5/3$ (see appendix Section A). For a self-similar, non intermittent flow, in the inertial range of scales (namely much smaller than the energy-injection scales and larger than the dissipative scales) the $n^{th}$ order structure function is expected to scale as $S_{n}(\tau)\propto\tau^{np}$, where $p$ is a phenomenological constant (Frisch, 1995). It is therefore straightforward to see that $K$ becomes independent of $\tau$. However, in the presence of intermittency, this linear scaling does not hold and the simplest intermittency model can be given as $S_{n}(\tau)\propto\tau^{np+q{(n)}}$, where $q{(n)}$ is a nonlinear correction accounting for the intermittent structures. For the kurtosis, this leads to a power-law scaling $K(\tau)\sim\tau^{-\kappa}$, with $\kappa=q{(4)}/2q{(2)}$. Such a scaling, universally observed in fluid turbulence, has recently been quantitatively described in the case of solar wind turbulence as well (Di Mare et al., 2019; Hernández et al., 2021; Sorriso-Valvo et al., 2023). In this work, we study the scaling properties of $K$ of the magnetic field fluctuations at different heliospheric distances.
Finally, the magnetic energy spectra and normalized cross-helicity co-spectra are defined as $PSD=\hat{B_{i}}^{\dagger}\hat{B_{i}}$ and $\hat{\sigma_{c}}=(\hat{b_{i}}^{\dagger}\hat{v_{i}}+\hat{v_{i}}^{\dagger}\hat{b_{i}})/(|\hat{b_{i}}|^{2}+|\hat{v_{i}}|^{2})$ respectively, where $\hat{B_{i}}$, $\hat{b_{i}}$ and $\hat{v_{i}}$ are the Fourier transforms of $B_{i}$, $b_{i}$ and $v_{i}$, respectively, with summation being intended over the repeated indices (where $i=r,t,n$).